George has 13.5 square yards of material, which comes in a width of 54 inches or 1.5 yards. How long is the rectangular piece of material?
Answer by edjones(8007)

1.5^2+b^2=13.5
b^2=12
b=sqrt(12)
=sqrt(4*3)
=2sqrt(3) yards
.
Ed